# ELFATT: Efficient Linear Fast Attention for Vision Transformers

### v1

@@ III-B General Attention Mechanism
III-C Kernelized Attention Mechanism
@@
$a=\phi\left(\textit{{x}}\right)\phi\left(\textit{{y}}\right)^{\top}.$ (3)
Performer [11] has proved that one of the best choices of the non-negative feature map $\phi$ for Eq. (3) is $\rm exp\left(\cdot\right)$. Performer obtains an exact alternative of Eq. (1) using random feature maps as follows,
${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)=\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)e\left(\textit{{x}}\right)e\left(\textit{{y}}\right)\right],$ (4)
@@
${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)=\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)e\left(\textit{{y}}\right)\right].$ (5)
Eq. (5) only holds when taking the sum of an infinite number of random vectors $\omega$. To avoid performing summation of infinite terms, Performer samples $c\times\log(c)$ random vectors $\omega$ to ensure a low approximation error. If $e\left(\textit{{y}}\right)$ is much smaller than ${\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)$, Eq. (5) can be further simplified and approximated as follows,
@@
${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)\approx\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)\right].$ (6)
Efficient attention (EFFATT) [18] is a special case of Eq. (6). It simplifies $\omega$ to a one-hot vector and only uses $c$ one-hot vectors in Eq. (6) to obtain an approximation as follows,
@@
${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)\approx\frac{1}{c}{\rm exp}\left(\textit{{x}}\right){\rm exp}\left(\textit{{y}}\right)^{\top}.$ (7)
@@
Its approximation error was studied and obtained in [5]. Eq. (7) has a problem of concentration reduction of attention maps [5] and Performer also has this problem when the number of $\omega$ sampled is too small. When the number of $\omega$ sampled is too large, Performer may be slower than the vanilla softmax-based attention mechanism.
III-D Efficient Linear Fast Attention Mechanism
@@
To address the problem of concentration reduction of Eq. (7), in this paper, we propose a novel attention mechanism as follows,
$\begin{split}{\rm exp}&\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\\ &\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)\right],\end{split}$ (8)

### v4

@@ 3.2. General Attention Mechanism
3.3. Kernelized Attention Mechanism
@@
(3) $a=\phi\left(\textit{{x}}\right)\phi\left(\textit{{y}}\right)^{\top}.$
Performer (Choromanski et al., 2021) has proved that one of the best choices of the non-negative feature map $\phi$ for Eq. (3) is $\rm exp\left(\cdot\right)$. Performer obtains an exact alternative of Eq. (1) using random feature maps as follows,
(4) ${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)=\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)e\left(\textit{{x}}\right)e\left(\textit{{y}}\right)\right],$
@@
(5) ${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)\approx\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)e\left(\textit{{y}}\right)\right].$
After the normalization, the left and right-hand sides of Eq. (5) will be equal. Eq. (4) only holds when taking the sum of an infinite number of random vectors $\omega$. To avoid performing summation of infinite terms, Performer samples $c\times\log(c)$ random vectors $\omega$ to ensure a low approximation error. If the change of $e\left(\textit{{y}}\right)$ is much smaller than ${\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)$, Eq. (5) can be further simplified and approximated as follows,
@@
(6) ${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)\approx\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)\right].$
Efficient attention (EFFATT) (Shen et al., 2021) is a special case of Eq. (6). It simplifies $\omega$ to a one-hot vector and only uses $c$ one-hot vectors in Eq. (6) to obtain an approximation as follows,
(7) ${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)\approx\frac{1}{c}{\rm exp}\left(\textit{{x}}\right){\rm exp}\left(\textit{{y}}\right)^{\top}.$
@@
Its approximation error was studied and obtained in (Wu et al., 2024). Eq. (7) has a problem of concentration reduction of attention maps (Wu et al., 2024) and Performer also has this problem when the number of $\omega$ sampled is too small. When the number of $\omega$ sampled is too large, Performer may be slower than VaniATT.
3.4. Efficient Linear Fast Attention Mechanism
@@
To address the defects of Eq. (7), in this paper, we propose a novel attention mechanism as follows,
(8) $\begin{split}{\rm exp}&\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)\right],\end{split}$